# Online Multivariate Changepoint Detection: Leveraging Links With Computational Geometry

### v2

@@ Appendix G An approximation for large $p$
In the section we consider the multivariate Gaussian case. That is we consider a sequence of $\{y_{t}\}_{t=1,\ldots}$ with elements in $\mathbb{R}^{p}$, with $y_{t}=\mu_{t}+\varepsilon_{t}$, and $\varepsilon_{t}$ in $\mathbb{R}^{p}$ i.i.d. $\mathcal{N}_{p}(0,I_{p}).$ We consider three types of statistics: a sparse statistics $S_{s,n}$, a dense statistics $D_{s,n}$ and a one-dimensional statistics $O_{s,n}^{i}$ defined as follow:
$S_{s,n}$ $=$ $(\max_{i}\lvert O_{s,n}^{i}\rvert)^{2}\,,$
$D_{s,n}$ $=$ $\biggl{\lVert}\frac{1}{\sqrt{(n-s)}}\sum_{t=s+1}^{n}\varepsilon_{t}\biggl{\rVert}_{2}^{2}\,,$
$O_{s,n}^{i}$ $=$ $\frac{1}{\sqrt{(n-s)}}\sum_{t=s+1}^{n}\varepsilon^{i}_{t}\,.$
where $\varepsilon^{i}_{t}$ is the $i-$th coordinate of $\varepsilon_{t}$ in $\mathbb{R}^{p}.$ In this appendix, using the arguably standard peeling arguments [65, 68] on all three statistics we demonstrate (1) that by using a time-varying threshold (i.e. increasing with the number of observations $n$), we can control the probability of detecting a change when there is none, that is the probability of false alarms, (2) that by using some fixed (independent of time) threshold we can control the average run length (ARL), and (3) that for a fixed threshold we can control the difference between the detection time $T$ and the true location of the change (that is the detection delay) in expectation and probability.
H.1 Probability of false alarms in the dense and sparse case
Corollary 2.
Considering a time-varying threshold of $b^{\text{sparse}}_{n,\alpha}=8\log(np/\alpha)$ for the sparse statistics and $b^{\text{dense}}_{n,\alpha}=p+2\sqrt{px_{n,\alpha}}+2x_{\alpha,n}$ with $x_{n,\alpha}=4\log(n)-log(\alpha)+1$ for the dense statistics we get that
@@ H.1 Probability of false alarms in the dense and sparse case
    Assuming there is no change, the probability of detecting a change is bounded by $2\alpha$. More precisely
    $\mathbb{P}(\exists s,n\in\mathbb{N}^{2}\ |\ S_{s,n}>\ b^{\text{sparse}}_{n,\alpha})\leq\alpha\quad\text{, and}\quad\mathbb{P}(\exists s,n\in\mathbb{N}^{2}\ |\ D_{s,n}>\ b^{\text{dense}}_{n,\alpha}))\leq\alpha\,.$
@@
    Assuming there is a change at $\tau$ the probability of detecting a change prior to $\tau$ is bounded by $2\alpha.$ More precisely
    $\mathbb{P}(\exists s,n\in\{1,\ldots,\tau\}^{2}\ |\ S_{s,n}>\ b^{\text{sparse}}_{n,\alpha})\leq\alpha\quad\text{, and}\quad\mathbb{P}(\exists s,n\in\{1,\ldots,\tau\}^{2}\ |\ D_{s,n}>\ b^{\text{dense}}_{n,\alpha}))\leq\alpha\,.$
@@
We follow step 1 and step 2 of the proof of Theorem 1 in [65] using peeling concentration bounds of our Lemma 4. ∎
@@
We can also control the average run length with a fixed threshold. Let’s assume we want an average run length of $\gamma$. Noting that $b^{\text{sparse}}_{n,\alpha}$ and $b^{\text{dense}}_{n,\alpha}$ are increasing with $n$, using two fixed thresholds $b^{\text{sparse}}_{n_{max},\alpha}$ and $b^{\text{dense}}_{n_{max},\alpha}$ we can lower bound the average run length by $(1-2\alpha)n_{max}.$ Thus taking $\alpha=1/4$ and $n_{max}=2\gamma$ we obtain the desired average-run-length.
H.2 Average detection delay for a fixed threshold
Assume that a change in mean occurs immediately after time $\tau$. Assume we are using a test statistic with a fixed threshold (following Remark 8) and let $T$ be the first time that our test statistic exceeds the threshold. The average detection delay is defined as
$\mbox{ADD}=\sup_{y_{1},\ldots,y_{\tau}}\mathbb{E}(T-\tau|T\geq\tau).$
That is, it is the worst-case expected time from the change until it is detected, conditional on no detection prior to $\tau$.
@@ H.2 Average detection delay for a fixed threshold
For a change of size $\delta$, with $\lVert\delta\rVert_{2}^{2}$ its square norm, the average detection delay for the dense statistic and threshold $c$ satisfies
$\mbox{ADD}\leq\frac{c+\lVert\delta\rVert_{2}\sqrt{8/\pi}}{\lVert\delta\rVert_{2}^{2}}+1.$
@@
Let $\tau$ be the time of the change, and $T$ the time at which the dense statistic first exceeds $c$. Consider the test statistic for a change at $\tau$ of size $\delta$, which at time $\tau+t$, for $t=1,2,\ldots$ is of the form:
$A_{t}=\sum_{i=1}^{t}\left(2\delta^{T}y_{\tau+i}-\lVert\delta\rVert_{2}^{2}\right).$ (23)
Define $A_{0}=0$. Now for any data $y_{1:t+\tau}$ we have $A_{t}\leq\max_{s=0}^{\tau+t-1}D_{s,\tau+t}$. Let $T^{\prime}$ be the time that $A_{t}$ first exceeds $c$. It follows that conditional on $T>\tau$, we have $T^{\prime}\geq T-\tau$. Thus we can bound the ADD by $\mathbb{E}(T^{\prime})$.
Denote each term in the sum on the right-hand side of (23) by $a_{i}$. We have these terms are independent, normally distributed with mean $\mathbb{E}(a_{i})=\delta^{T}\delta=\lVert\delta\rVert_{2}^{2}$ and variance $4\lVert\delta\rVert_{2}^{2}$. Let $\mu=\lVert\delta\rVert_{2}^{2}$. For each $t$ it is trivial to see that $\mathbb{E}(|A_{t}|)<\infty$, and thus $A_{t}-t\mu$ is a Martingale. It is also straightforward to show that $T^{\prime}$ is a stopping time and $\mathbb{E}(T^{\prime})<\infty$ so we can apply the optional stopping theorem. This gives
$\mathbb{E}\left(A_{T^{\prime}}-T^{\prime}\mu\right)=0.$
Thus $\mathbb{E}(T^{\prime})=\mathbb{E}(A_{T^{\prime}})/\mu$. We can bound $\mathbb{E}(A_{T^{\prime}})$ from above by considering the distribution of $A_{T^{\prime}}$ conditional on $A_{T^{\prime}-1}$. We have
$\mathbb{E}(A_{T^{\prime}}|A_{T^{\prime}-1}=A)=\mathbb{E}(A+\mu+\sqrt{4\mu}\epsilon|A+\mu+\sqrt{4\mu}\epsilon\geq c),$
@@
$\mathbb{E}(A_{T^{\prime}})\leq\mathbb{E}(c+\mu+\sqrt{2\mu}|\epsilon|)=c+\mu+\sqrt{2\mu}\sqrt{\frac{2}{\pi}}.$
@@
It is simple to extend this result to get a bound on the average detection delay of the sparse statistic; as the same argument applies to the statistic that considers any given variate changing. $\lVert\delta\rVert_{2}$ would simplify to $|\delta_{j}|$ with $j$ the variate that is changing.
H.2.1 Probabilistic bound on the Detection Delay in the dense case
The bound on the ADD in Lemma 3 is proportional to the threshold and therefore in the dense case (see Lemma 2) it includes a linear term in $p.$ We argue that this is likely sub-optimal due to our proof technique. Considering the dense statistics and following Remark 8 we control the ARL at $\gamma$ taking a fixed threshold of $b_{2\gamma,.25}=p+2\sqrt{px_{2\gamma,.25}}+2x_{2\gamma,.25}$ with $x_{2\gamma,.25}=4\log(2\gamma)-log(\alpha)+1.$ Now using Lemma 5 and a maximum run length of $2\gamma$ we can further control the detection delay in probability. Assume there is a change at $\tau$ of $\delta$ we have for any $n>\tau:$
$max_{s,n}D_{s,n}$$\geq$ $\biggl{\lVert}\sum_{t=\tau+1}^{n}\frac{y_{t}}{\sqrt{n-\tau}}\biggl{\rVert}_{2}^{2}=\biggl{\lVert}\sum_{t=\tau+1}^{n}\frac{\varepsilon_{t}}{\sqrt{n-\tau}}+\sqrt{n-\tau}\delta\biggl{\rVert}_{2}^{2}$
$\geq$ $\lVert\delta\rVert_{2}^{2}(n-\tau)+2\sqrt{n-\tau}\frac{\sum_{\tau+1}^{n}\langle\delta,\varepsilon_{i}\rangle}{\lVert\delta\rVert\sqrt{n-\tau}}+\frac{\lVert\sum_{\tau+1}^{n}\varepsilon_{i}\rVert_{2}^{2}}{{n-\tau}}\,.$
Notice that $\frac{\langle\delta,\varepsilon_{i}\rangle}{\lVert\delta\rVert}$ are i.i.d. Gaussian. We can thus apply Lemma 5. We can also use Lemma 4 Equation (26) on $\frac{\lVert\sum_{\tau+1}^{n}\varepsilon_{i}\rVert_{2}^{2}}{{n-\tau}}.$ Now, considering a maximum run length of $2\gamma$, we get that with probability at least $1-2\alpha$
$max_{s,n}D_{s,n}$$\geq$ $\lVert\delta\rVert_{2}^{2}(n-\tau)-4\lVert\delta\rVert\sqrt{(n-\tau)x^{\prime}_{2\gamma,\alpha}}+p-2\sqrt{x_{2\gamma,\alpha}}$
with $x^{\prime}_{2\gamma,\alpha}=2\log(\log(6\gamma))-\log(\alpha)+1.$ Considering the previous equation as a second degree polynomial in $\sqrt{n-\tau}\lVert\delta\rVert$ and searching for the first $n^{*}$ such that it is superior to $b_{2\gamma,.25}$ we get that $n^{*}$ is the smallest integer such that
$\lVert\delta\rVert_{2}\sqrt{n-\tau}$$\geq$ $2\sqrt{x^{\prime}_{2\gamma,\alpha}}+\sqrt{4x^{\prime}_{2\gamma,\alpha}+2\sqrt{px_{2\gamma,\alpha}}+2\sqrt{px_{2\gamma,.25}}+2x_{2\gamma,.25}}.$
Taking the square and using the fact that $(a+b)^{2}\leq 2a^{2}+2b^{2}$ we get with probability at least $1-2\alpha$ that the detection time is such that
$T-\tau$$\leq$ $\frac{1}{\lVert\delta\rVert_{2}^{2}}\left(16x^{\prime}_{2\gamma,\alpha}+4\sqrt{px_{2\gamma,\alpha}}+4\sqrt{px_{2\gamma,.25}}+4x_{2\gamma,.25}\right)+1\,.$
Observe that numerator is in $\lVert\delta\rVert_{2}^{2}$ and the denominator has no linear term in $p.$
H.3 Some peeling bounds
In this subsection we provide bounds on the sparse ($S_{s,n}$) and dense ($D_{s,n}$) statistics, which are useful to control the probability of false alarms (see Corollary 2). We then recall a bound by [68] which is useful to study the detection delay.

### v3

@@ I.1 Definitions and problem statement
In this section, we consider the multivariate Gaussian case. That is, we consider a sequence $\{y_{t}\}_{t=1,2,\ldots}$ with elements in $\mathbb{R}^{p}$, with $y_{t}=\mu_{t}+\varepsilon_{t}$, and $\varepsilon_{t}$ i.i.d $\mathcal{N}_{p}(0,I_{p})$. At time $n$, for any possible change $\tau<n$, we consider two types of statistics:
  • Ranked-based statistics, $S^{\#s}_{\tau,n}$, considering the $s$ dimensions with the most evidence for a change at $\tau$;
  • Threshold-based statistics, $S^{(a)}_{\tau,n}$, considering, as in Chen et al. [2022], only the dimensions for which the evidence for a change exceeds the threshold $a$.
To formally present these statistics, we first define the unsorted per-dimension evidence for a change at $\tau$ at time $n$ knowing the pre-change mean to be 0, $O^{i}_{\tau,n}(0)$ or not knowing the pre-change mean, $O^{i}_{\tau,n}(.)$:
$O^{i}_{\tau,n}(0)=\frac{1}{\sqrt{n-\tau}}\sum_{t=\tau+1}^{n}\varepsilon^{i}_{t},\qquad O^{i}_{\tau,n}(.)=\sqrt{\frac{\tau(n-\tau)}{n}}\left(\frac{1}{n-\tau}\sum_{t=\tau+1}^{n}\varepsilon^{i}_{t}-\frac{1}{\tau}\sum_{t=1}^{\tau}\varepsilon^{i}_{t}\right),$
where $\varepsilon^{i}_{t}$ is the $i-$th coordinate of $\varepsilon_{t}$ in $\mathbb{R}^{p}$. We also define the sorted per-dimension evidence for a change at $\tau$ at time $n$, $O^{\#j}_{\tau,n}(0)$ as the $j$-th highest value out of all $O^{i}_{\tau,n}(0)$. Similarly, for the unknown pre-change mean scenario, we define $O^{\#j}_{\tau,n}(.)$ as the $j$-th highest value out of all $O^{i}_{\tau,n}(.)$. Based on that, we now define our rank and threshold-based statistics knowing the pre-change mean to be 0 as:
$S^{\#s}_{\tau,n}(0)=\sum_{j=1}^{s}(O^{\#i}_{\tau,n}(0))^{2}\quad\text{and}\quad S^{(a)}_{\tau,n}(0)=\sum_{i=1}^{p}(O^{i}_{\tau,n}(0))^{2}\mathbbm{1}_{|O^{i}_{\tau,n}(0)|\geq a}\,,$ (19)
where $\mathbbm{1}$ denotes the indicator function. We define similar statistics if the pre-change mean is not known:
$S^{\#s}_{\tau,n}(.)=\sum_{j=1}^{s}(O^{\#i}_{\tau,n}(.))^{2}\quad\text{and}\quad S^{(a)}_{\tau,n}(.)=\sum_{i=1}^{p}(O^{i}_{\tau,n}(.))^{2}\mathbbm{1}_{|O^{i}_{\tau,n}(.)|\geq a}\,.$ (20)
For simplicity, when this does not cause confusion, we will omit the $(0)$ and (.)(.)( . ) notations. It will sometimes make sense to consider separately the ranked-1 statistics $S^{\#1}_{\tau,n}$. We will refer to $S^{\#p}_{m,n}=S^{(a=0)}_{m,n}$ as the dense statistic.
First we summarise our results:
@@
    For all ranked and thresholded statistics, we provide, in Section I.2, time-varying thresholds to control their false alarm rate. To be specific, considering a generic ranked or thresholded statistic $\tilde{S}_{\tau,n}$, we provide a varying threshold $c_{n}(\alpha)$ for which:
    $\mathbb{P}(\exists\ \tau<n\in\mathbb{N}\ \text{such that }\,\tilde{S}_{\tau,n}\geq c_{n}(\alpha))\ \leq\ \alpha.$
    Importantly, our thresholds are also valid if the pre-change mean is unknown. As a byproduct, we can calibrate a fixed threshold to control the Average Run Length, $\gamma$. That is, we can define a threshold $c(\gamma)$ such that if all $\mu_{t}$ are equal to 00 then $E(\tilde{T})\geq\gamma,$ where $\tilde{T}$ is the first time our generic statistic passes the threshold.
@@
    Second, in Section I.3, assuming the pre-change mean is known, and for a fixed threshold $c$, we control using the optional stopping time theorem, the Average Detection Delay (ADD). To be specific, assuming $\mu_{t}=0$ for all $t\leq\tau^{*}$ and $\mu_{t}=\delta$ otherwise we get a bound on
    $\sup_{y_{1},\ldots,y_{\tau^{*}}}\mathbb{E}(\tilde{T}-\tau^{*}|\tilde{T}\geq\tau^{*}),$
    involving only $\lVert\delta\rVert$ and the threshold. For thresholded statistics, the same result holds by replacing $c$ by $c+pa^{2}$ (see remark 6).
  3. 3.
    Finally, in Section I.4, for all our ranked and thresholded statistics, assuming we know the pre-change mean, we provide probabilistic bounds on the Detection Delay. To be specific, again assuming $\mu_{t}=0$ for all $t\leq\tau^{*}$ and $\mu_{t}=\delta$ otherwise, we bound with probability $1-\alpha$ the detection time, $\tilde{T}-\tau^{*}$, by a term involving only $\lVert\delta\rVert$, the threshold, and $\alpha$.
I.2 Probability of false alarms for all statistics
Proposition 8.
For a sequence without change : $y_{t}=\varepsilon_{t}$ with $\varepsilon_{t}$ i.i.d. $\mathcal{N}_{p}(0,I_{p})$. Considering a time-varying $x_{n}(\alpha)=4\log(n)-\log(\alpha),$ we can control the false alarm rate of ranked and thresholded statistics at level $\alpha$. That is for a generic statistic $\tilde{S}_{\tau,n}$ we control:
$\mathbb{P}(\exists\ \tau<n\in\mathbb{N}\ \text{such that }\,\tilde{S}_{\tau,n}\geq c_{n}(\alpha))\ \leq\ \alpha,$
using the thresholds in table 10.
Statistic $\tilde{S}_{\tau,n}$ Threshold $c_{n}(\alpha)$
$S^{\#1}_{\tau,n}$ $2x_{n}(\alpha)+2\log(2p)$
$S^{\#s}_{\tau,n}$ $2(x_{n}(\alpha)+\log{p\choose s})+2\sqrt{s(x_{n}(\alpha)+\log{p\choose s})}+s$
$S^{(a)}_{\tau,n}$ $4x_{n}(\alpha)+6pe^{-a^{2}/8}$
Table 10: Probability bounds for our statistics with $x_{n}(\alpha)=4\log(n)-\log(\alpha)$
@@ I.2 Probability of false alarms for all statistics
We follow Step 1 and Step 2 of the proof of Theorem 1 in Yu et al. [2023] using the peeling concentration bounds of our Lemma 7. ∎
Importantly, when using MdFOCuS, one typically considers several statistics. Denoting #$\#Stats$ the number of statistics, we control simultaneously all of them at level $\alpha^{\prime}$ taking $\alpha=\alpha^{\prime}/\#Stats$ in the thresholds of Table 10.
@@
Using Proposition 8 we can also control the Average Run Length (ARL) using a fixed threshold, $c$. Consider the Rank-1 statistic, $S^{\#1}_{\tau,n}$, as an example. We note that its threshold $2x_{n}(\alpha)+2\log(2p)$ increases with $n$. Thus, using a fixed threshold $c=2x_{N}(\alpha)+2\log(p)$ for all time $n$, the probability of stopping before $N$ is less than $\alpha$, and therefore the ARL is larger than $(1-\alpha)N$. To control the ARL at any desired level $\gamma$, it suffices to choose a sufficiently large $N$ and a sufficiently small probability $\alpha$ such that $(1-\alpha)N\geq\gamma$.
I.3 Average detection delay for a fixed threshold and knowing the pre-change mean
In this section, we assume that a change in mean occurs immediately after time $\tau^{*}$. Assume that we are using the dense statistic $S^{\#p}_{\tau,n}(0)=S^{(0)}_{\tau,n}(0)$ with a fixed threshold (following Remark 4) and let $T$ be the first time that our test statistic exceeds the threshold. The average detection delay is defined as
$\mbox{ADD}=\sup_{y_{1},\ldots,y_{\tau^{*}}}\mathbb{E}(T-\tau^{*}|T\geq\tau^{*}).$
That is, it is the worst-case expected time from the change until it is detected, conditional on no detection before $\tau^{*}$.
@@ I.3 Average detection delay for a fixed threshold and knowing the pre-change mean
For a change of size $\delta$, with $\lVert\delta\rVert^{2}$ its Euclidean squared norm, the average detection delay for the dense statistic with a known pre-change mean ($S^{\#p}_{\tau,n}(0)=S^{(0)}_{\tau,n}(0)$) and a fixed threshold $c$ satisfies
$\mbox{ADD}\leq\frac{c+\lVert\delta\rVert\sqrt{8/\pi}}{\lVert\delta\rVert^{2}}+1.$
@@
Let $\tau^{*}$ be the time of the change, and $T$ the time at which the dense statistic first exceeds $c$. Consider the test statistic for a change at $\tau$ of size $\delta$, which at time $\tau^{*}+t$, for $t=1,2,\ldots$ is of the form:
$A_{t}=\sum_{i=1}^{t}\left(2\delta^{T}y_{\tau^{*}+i}-\lVert\delta\rVert^{2}\right).$ (21)
Note that $\max_{\delta}A_{t}=S^{\#p}_{\tau^{*},\tau^{*}+t}$. Define $A_{0}=0$.
Now for any data $y_{1:t+\tau^{*}}$ we have $A_{t}\leq\max_{\tau=0}^{\tau+t-1}S^{\#p}_{\tau,t+\tau^{*}}(0)$. Let $T^{\prime}$ be the time that $A_{t}$ first exceeds $c$. It follows that conditional on $T>\tau^{*}$, we have $T^{\prime}\geq T-\tau^{*}$. Thus we can bound the ADD by $\mathbb{E}(T^{\prime})$.
Denote each term in the sum on the right-hand side of (21) by $a_{i}$. We have these terms are independent, normally distributed with mean $\mathbb{E}(a_{i})=\delta^{T}\delta=\lVert\delta\rVert^{2}$ and variance $4\lVert\delta\rVert^{2}$. For each $t$ it is trivial to see that $\mathbb{E}(|A_{t}|)<\infty$, and thus $A_{t}-t\lVert\delta\rVert^{2}$ is a Martingale. It is also straightforward to show that $T^{\prime}$ is a stopping time and $\mathbb{E}(T^{\prime})<\infty$ so we can apply the optional stopping theorem. This gives
$\mathbb{E}\left(A_{T^{\prime}}-T^{\prime}\lVert\delta\rVert^{2}\right)=0.$
Thus $\mathbb{E}(T^{\prime})=\mathbb{E}(A_{T^{\prime}})/\lVert\delta\rVert^{2}$. We can bound $\mathbb{E}(A_{T^{\prime}})$ from above by considering the distribution of $A_{T^{\prime}}$ conditional on $A_{T^{\prime}-1}$. We have
$\mathbb{E}(A_{T^{\prime}}|A_{T^{\prime}-1}=A)=\mathbb{E}\left(A+\lVert\delta\rVert^{2}+\sqrt{4\lVert\delta\rVert^{2}}\epsilon\ |\ A+\lVert\delta\rVert^{2}+\sqrt{4\lVert\delta\rVert^{2}}\epsilon\geq c\right),$
@@
$\mathbb{E}(A_{T^{\prime}})\leq\mathbb{E}(c+\lVert\delta\rVert^{2}+\sqrt{2\lVert\delta\rVert^{2}}|\epsilon|)=c+\lVert\delta\rVert^{2}+\sqrt{2\lVert\delta\rVert^{2}}\sqrt{\frac{2}{\pi}}.$
@@
It is simple to extend this result to get a bound on the ADD of any ranked statistics $S^{\#s}_{\tau,n}(0)$, because the same argument applies to any statistic that considers a given subset of $s$ variates changing. Following the notation introduced in Appendix G (on the approximation), denoting $R$ the subset of $s$ variates that are changing, we would recover the same formula considering $y^{[R]}$ in $\mathbb{R}^{s}$ and replacing $\lVert\delta\rVert$ by $\lVert\delta^{[R]}\rVert$.
Remark 6.
The proof also applies to any thresholded statistics with threshold $a$. We replace $c$ by $c+pa^{2}$ in the proof, note that $S^{\#p}_{\tau,n}-pa^{2}$ is always smaller than $S^{(a)}_{\tau,n}$ and the bound follows.
I.4 Probabilistic bound on the Detection Delay for a fixed threshold and knowing the pre-change mean
In this section, we again assume that a change in mean occurs immediately after time $\tau^{*}$, calling $T$ the first time our statistic passes the threshold $c_{n}$ we give probabilistic bounds on the detection delay $T-\tau^{*}$. We start with ranked statistics and then proceed with thresholded statistics.
Proposition 9.
Consider a change of size $\delta$ at time $\tau^{*}$, that is $y_{t}=\varepsilon_{t}$ for $t\leq\tau^{*}$ and $y_{t}=\delta+\varepsilon_{t}$ for $t>\tau^{*}$ with $\varepsilon_{t}\sim\mathcal{N}(0,I_{p})$ and $\lVert\delta\rVert>0$. We denote $T$ the first time the dense statistic, $S^{\#p}_{\tau,n}(0)$, passes the fixed positive threshold $c$. With probability at least $(1-\alpha)$ the detection delay, $T-\tau^{*}$, is less than
$\frac{2c-8\log(\alpha/2)}{\lVert\delta\rVert^{2}}$
In addition, if $c>p$, then with probability at least $(1-\alpha)$,
$T-\tau^{*}<\frac{2(c-p)+2\sqrt{p\log(3/\alpha)}-8\log(\alpha/3)}{\lVert\delta\rVert^{2}}$
Proof.
We consider the statistic $S^{\#p}_{\tau^{*},n}(0)$ for increasing $n$ and apply Lemma 8 and find that for any time larger than the provided bound, the statistic would pass the threshold with probability at least $(1-\alpha)$. Hence, the detection delay would be smaller with probability $(1-\alpha)$. ∎
Remark 7.
It is simple to extend this result to get a bound on the detection delay of any ranked statistic $S^{\#s}_{\tau,n}(0)$. This is because the same argument applies to the statistic that considers any subset of $s$ variates changing. Following the notation introduced in Appendix G (on the approximation), denoting $R$ the subset of $s$ variates that are changing, we would consider $y^{R}$ in $\mathbb{R}^{s}$ and recover the same formula replacing $\lVert\delta\rVert$ by $\lVert\delta^{[R]}\rVert$.
Using Lemma 8 we could get similar bounds for a time-varying threshold assuming it is positive, increasing, and concave thresholds.
Before providing bounds on the detection delay of thresholded statistics, we recall the notion of effective sparsity defined in Chen et al. [2022]. For any $\delta$ in $\mathbb{R}^{p}$ such that $\lVert\delta\rVert>0$ the effective sparsity $z(\delta)$ is the smallest integer in $\{2^{0},2^{1},\ldots,2^{\lfloor\log_{2}(p)}\rfloor\}$ such that the set
$\mathcal{Z}(\delta)=\{j\ \text{such that}\ |\delta^{j}|>\lVert\delta\rVert/(z(\delta)\log_{2}(p))\},$
has cardinality at least $z(\delta)$.
Proposition 10.
Consider a change of size $\delta$ at time $\tau^{*}$, that is $y_{t}=\varepsilon_{t}$ for $t\leq\tau^{*}$ and $y_{t}=\delta+\varepsilon_{t}$ for $t>\tau^{*}$ with $\varepsilon_{t}\sim\mathcal{N}(0,I_{p})$ and $\lVert\delta\rVert>0$. We denote $T$ the first time the thresholded statistic, $S^{(a)}_{\tau,n}(0)$, passes the fixed threshold $c$. With probability at least $(1-\alpha)$ the detection delay, $T-\tau^{*}$, is less than than
$\log_{2}(p)\max\left\{\frac{2c-8\log(\alpha)+8\log(2(z(\delta)+1)))}{\lVert\delta\rVert^{2}},\frac{2a^{2}-8\log(\alpha)+8\log(2(z(\delta)+1)))}{\lVert\delta\rVert^{2}/z(\delta)}\right\}$
Proof.
We consider a subset $\mathcal{Z^{\prime}}$ of size $z(\delta)$ of $\mathcal{Z}(\delta)$. All variates $j$ in this subset validate the effective sparsity constraint: $|\delta_{j}|>\lVert\delta\rVert/(z(\delta)\log_{2}(p))$. We first lower bound the contribution to $S^{(a)}_{\tau,n}$ of all variates not in $\mathcal{Z^{\prime}}$ by 00. Then we consider a time large enough to ensure that simultaneously
  • the contribution of each variates in $\mathcal{Z^{\prime}}$ passes the threshold $a$
  • the sum of these $z(\delta)$ contributions passes the fixed threshold $c$.
To this end we use Lemma 8 exactly $z(\delta)+1$ times:
  • for each variates in $\mathcal{Z^{\prime}}$ the lemma applies replacing $\lVert\delta\rVert^{2}$ by $\frac{\lVert\delta\rVert^{2}}{z(\delta)\log_{2}(p)}$ and $\alpha$ by $\frac{\alpha}{z(\delta)+1}$ (in preparation for a union bound).
  • for the sum the lemma applies replacing $\lVert\delta\rVert^{2}$ by $\frac{\lVert\delta\rVert^{2}}{\log_{2}(p)}$ and $\alpha$ by $\frac{\alpha}{z(\delta)+1}$. This is because, by definition of $z(\delta)$ the squared euclidean norm restricted to the variates in $\mathcal{Z^{\prime}}$ is such that $\lVert\delta^{[\mathcal{Z^{\prime}}]}\rVert^{2}\geq\frac{\lVert\delta\rVert^{2}}{\log_{2}(p)}$.
We get the desired result on the detection delay by taking a union bound over these $z(\delta)+1$ events. ∎
Using Lemma 8 we could get similar bounds for a time-varying threshold assuming it is positive, increasing, and concave thresholds.
I.5 Some concentration bounds
In this subsection, we provide concentration bounds on our ranked and threshold-based statistics, respectively $S^{\#s}_{\tau,n}$ and $S^{(a)}_{\tau,n}$ defined in (19) and (20). These bounds are useful to control the run-length and the detection delay
First, exploiting the peeling argument of Yu et al. [2023] we provide bounds considering all possible values of $\tau$ and $n$ which are useful to control the probability of false alarms (see Proposition 8). These bounds and their proof are valid for both the known and unknown pre-change mean scenarios. This is because if there is no change $O^{i}_{\tau,n}(0)$ and $O^{i}_{\tau,n}(.)$ are both Gaussian with mean zero and variance one.